OpenStudy (anonymous):

Find the real zeros of the trigonometric function on the interval 0 ≤ x < 2π f(x)=2cos^2x+2sinx-3

OpenStudy (solomonzelman):

use \(\large\color{black}{ \bf sin^2x+cos^2x=1 }\)

Parth (parthkohli):

\[f(x) = 2(1 - \sin^2 x) + 2\sin x - 3\]This is a quadratic in \(\sin(x).\) For easy work, let \(\sin(x) = t\) and find the roots.

OpenStudy (anonymous):

So would it be sin = -1/2 and i look for those points on the unit circle>

Parth (parthkohli):

No, I actually seem to be getting no values of sin from this.

OpenStudy (anonymous):

Im confused on how you got that, don't you distribute the 2 in front of the parenthesis and then set the equation equal to zero?

Parth (parthkohli):

Yes. I'm still getting no real solutions.

OpenStudy (anonymous):

Apparently I'm not doing something right then, this is how I've been solving the equation and after that it doenst make sense.